If the molar mass ratio between reactants A and B in a reaction is 4A:3B, how many moles of A will be needed to react with 4.5 mol of B?
1
Expert's answer
2019-05-21T05:54:54-0400

Regardless of molar mass of reactants, right wording of the question should be : "If the molar ratio between reactants A and B in a reaction is 4A:3B, how many moles of A will be needed to react with 4.5 mol of B?"


From stoichiometry of the reaction ratio A:B is 4:3. On the other hand A:B as x:4.5, where x is unknown amount of A reacted with 4.5 mol of B. From the equation we can find x:

"\frac{4}{3}=\frac{x}{4.5}"

"x=\frac{4}{3}\cdot 4.5 mol = 6 mol"

The answer: 6 moles of A will be needed to react with 4.5 mol of B.